OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)?

OpenStudy (anonymous):

y – 5 = –1(x – 2)

OpenStudy (aum):

First find the slope of the line from the given two points. Slope \(\Large m = \frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 0}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 8}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)